OpenStudy (anonymous):

how to prove that (1/(f(D_x,D_y))*e^(ax+by))=(1/f(a,b))*)e^(ax+by)? if q(x,y) exponential.

OpenStudy (polaris_s0i):

I assume you're question is: \[\frac{1}{f(Dx, Dy)} * e^(ax+by) = \frac{ 1 }{ f(a,b)} * e^(ax + by)\] The equation is rather vague, but if you take e^(ax+by) alone, its derivative is \[d(e^(ax+by)) = (adx + bdy)e^(ax+by)\] The q(x,y) isn't even in you're question. Is there more to this?

OpenStudy (anonymous):

q(x,y)=e^ax+by

OpenStudy (anonymous):

youre' right polaris,how to prove\[\frac{ 1 }{ f(Dx,Dy) }*e ^{ax+by}=\frac{ 1 }{ f(a,b) }*e ^{ax+by}\]

OpenStudy (polaris_s0i):

so is the Dx = Dx(q) and Dy = Dy(q)?

OpenStudy (polaris_s0i):

because you equation is correct if (ax + by) is the function you are taking the derivative of.

OpenStudy (anonymous):

derivative of?

OpenStudy (polaris_s0i):

Dx is partial Derivative of X correct? Partial of (ax + by) with respect to X is a partial of (ax + by) with respect to Y is b, Or am I way off of what this question is asking?

OpenStudy (polaris_s0i):

been learning more latex so I can type this properly: \[\frac{1}{f(\partial_x, \partial_y)}e^{ax+by} = \frac{1}{f(a,b)}e^{ax+by}\] so what I'm saying is that if \[P = (ax + by)\] \[P_x = \frac{\partial}{\partial x}(ax+by) = a\] \[P_y = \frac{\partial}{\partial y}(ax+by) = b\] \[f(P_x,P_y) = f(a,b)\] I'm just not sure this is what you are looking for or not, because it seems way to obvious.

OpenStudy (anonymous):

i think you has been help me